• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name glashausser

    Log entry time 16:12:46 on March24,2000

    Entry number 36829

    keyword=shift summary

    8:00 Running Kin#31 at 105 microA
    9:00 Beam off for energy studies, then optics by MCC
    9:30 Eric V to MCC for Harp scan, try to put in dummy target, seems stuck
    9:45 MCC makes controlled access for beam line quad power supply
    10:00 Rebooted target software and target moved
    10:55 still in controlled access
    11:10 power permit
    13:25 beam back for Kin#31, have to adjust raster--with old setting was 3x4.5
    (Run 2915 is junk, as is 2916)
    14:00 finally get raster right to get spot++ size of 3.5x3.5mm, [In "set magnet
    current" box on screen, "Raster box size" reads 2.0 x 4.0 mm]
    15:00 2917 stopped on its own for no good reason, noticed after 20 minutes
    restarted CODA 2x
    15:37 started run 2918
    16:00 energy lock had been with us for about the last hour, maybe more. Now back
    to Hall C